7.2 aliens = 1 monster
1 monster =15.5 oranges
Using the conversion above, how many oranges are equal to 1 alien?
A)0.46
B)1.95
C)2.15
D)22.7

Found 2 solutions by rothauserc, math_helper:
Answer by rothauserc(4718)

1 Alien = 1/7.2 = 0.1389 Monster
:
1 Alien = 0.1389 * 15.5 = 2.15295 is approximately 2.15 oranges
:
Answer is C. 2.15

Answer by math_helper(2461)


Using unit-analysis:

1 alien
= 1 alien * (1 monster / 7.2 aliens) * (15.5 oranges / 1 monster)
= 2.15278 oranges